OpenStudy (anonymous):

(4x+1)^2=20 solve each equation by the square root property

hartnn (hartnn):

did you take square root on both sides, first ? what you get ?

OpenStudy (anonymous):

solve for X i mean

hartnn (hartnn):

yeah, take square root on both sides of \((4x+1)^2=20\) what do you get ?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

4x-1=\[\sqrt{20}\] =

hartnn (hartnn):

that would be \(4x+1 =\pm \sqrt {20}\) ok? now subtract 1 from both sides, what u get ?

OpenStudy (anonymous):

4x=\[\pm \sqrt{20}\] - 1

hartnn (hartnn):

thats correct, now just divide both sides by 4 and you get 2 values of x.

OpenStudy (anonymous):

i got x = PM square root 4

hartnn (hartnn):

? how ? \(4x=\pm \sqrt{20}-1 \implies x = \dfrac{\pm \sqrt{20}-1}{4}\)

OpenStudy (anonymous):